$SPAD/src/input sregset.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool sregset.output
)set message test on
)set message auto off
)clear all
--S 1 of 23
R := Integer
--R
--R
--R   (1)  Integer
--R                                                                 Type: Domain
--E 1
--S 2 of 23
ls : List Symbol := [x,y,z,t]
--R
--R
--R   (2)  [x,y,z,t]
--R                                                           Type: List(Symbol)
--E 2
--S 3 of 23
V := OVAR(ls)
--R
--R
--R   (3)  OrderedVariableList([x,y,z,t])
--R                                                                 Type: Domain
--E 3
--S 4 of 23
E := IndexedExponents V
--R
--R
--R   (4)  IndexedExponents(OrderedVariableList([x,y,z,t]))
--R                                                                 Type: Domain
--E 4
--S 5 of 23
P := NSMP(R, V)
--R

--R
--R   (5)
--R   NewSparseMultivariatePolynomial(Integer,OrderedVariableList([x,y,z,t]))
--R                                                                 Type: Domain
--E 5
--S 6 of 23
x: P := 'x
--R
--R
--R   (6)  x
--RType: NewSparseMultivariatePolynomial(Integer,OrderedVariableList([x,y,z,t]))
--E 6
--S 7 of 23
y: P := 'y
--R
--R
--R   (7)  y
--RType: NewSparseMultivariatePolynomial(Integer,OrderedVariableList([x,y,z,t]))
--E 7
--S 8 of 23
z: P := 'z
--R
--R
--R   (8)  z
--RType: NewSparseMultivariatePolynomial(Integer,OrderedVariableList([x,y,z,t]))
--E 8
--S 9 of 23
t: P := 't
--R
--R
--R   (9)  t
--RType: NewSparseMultivariatePolynomial(Integer,OrderedVariableList([x,y,z,t]))
--E 9
--S 10 of 23
ST := SREGSET(R,E,V,P)
--R
--R
--R   (10)
--R  SquareFreeRegularTriangularSet(Integer,IndexedExponents(OrderedVariableList([
--R  x,y,z,t])),OrderedVariableList([x,y,z,t]),NewSparseMultivariatePolynomial(Int
--R  eger,OrderedVariableList([x,y,z,t])))
--R                                                                 Type: Domain
--E 10
--S 11 of 23

p1 := x ** 31 - x ** 6 - x - y
--R
--R
--R          31    6
--R   (11)  x   - x  - x - y
--RType: NewSparseMultivariatePolynomial(Integer,OrderedVariableList([x,y,z,t]))
--E 11
--S 12 of 23
p2 := x ** 8  - z
--R
--R
--R          8
--R   (12)  x  - z
--RType: NewSparseMultivariatePolynomial(Integer,OrderedVariableList([x,y,z,t]))
--E 12
--S 13 of 23
p3 := x ** 10 - t
--R
--R
--R          10
--R   (13)  x   - t
--RType: NewSparseMultivariatePolynomial(Integer,OrderedVariableList([x,y,z,t]))
--E 13
--S 14 of 23
lp := [p1, p2, p3]
--R
--R
--R           31    6          8      10
--R   (14)  [x   - x  - x - y,x  - z,x   - t]
--RType: List(NewSparseMultivariatePolynomial(Integer,OrderedVariableList([x,y,z,t])))
--E 14
--S 15 of 23
zeroSetSplit(lp)$ST
--R
--R
--R            5    4      2     3     8     5    3    2   4                2
--R   (15)  [{z  - t ,t z y  + 2z y - t  + 2t  + t  - t ,(t  - t)x - t y - z }]
--IType: List(SquareFreeRegularTriangularSet(Integer,...
--E 15
--S 16 of 23
zeroSetSplit(lp,false)$ST
--R
--R
--R   (16)
--R      5    4      2     3     8     5    3    2   4                2

--R   [{z  - t ,t z y  + 2z y - t  + 2t  + t  - t ,(t  - t)x - t y - z },
--R      3      5            2    2
--R    {t  - 1,z  - t,t y + z ,z x  - t}, {t,z,y,x}]
--IType: List(SquareFreeRegularTriangularSet(Integer,...
--E 16
--S 17 of 23
T := REGSET(R,E,V,P)
--R
--R
--R   (17)
--R  RegularTriangularSet(Integer,IndexedExponents(OrderedVariableList([x,y,z,t]))
--R  ,OrderedVariableList([x,y,z,t]),NewSparseMultivariatePolynomial(Integer,Order
--R  edVariableList([x,y,z,t])))
--R                                                                 Type: Domain
--E 17
--S 18 of 23
lts := zeroSetSplit(lp,false)$T
--R
--R
--R   (18)
--R      5    4      2     3     8     5    3    2   4                2
--R   [{z  - t ,t z y  + 2z y - t  + 2t  + t  - t ,(t  - t)x - t y - z },
--R      3      5          2     3         2
--R    {t  - 1,z  - t,t z y  + 2z y + 1,z x  - t}, {t,z,y,x}]
--IType: List(RegularTriangularSet(Integer,...
--E 18
--S 19 of 23
ts := lts.2
--R
--R
--R           3      5          2     3         2
--R   (19)  {t  - 1,z  - t,t z y  + 2z y + 1,z x  - t}
--IType: RegularTriangularSet(Integer,...
--E 19
--S 20 of 23
pol := select(ts,'y)$T
--R
--R
--R              2     3
--R   (20)  t z y  + 2z y + 1
--IType: Union(NewSparseMultivariatePolynomial(Integer,...
--E 20
--S 21 of 23
tower := collectUnder(ts,'y)$T
--R

--R
--R           3      5
--R   (21)  {t  - 1,z  - t}
--IType: RegularTriangularSet(Integer,IndexedExponents(...
--E 21
--S 22 of 23
pack := RegularTriangularSetGcdPackage(R,E,V,P,T)
--R
--R
--R   (22)
--R  RegularTriangularSetGcdPackage(Integer,IndexedExponents(OrderedVariableList([
--R  x,y,z,t])),OrderedVariableList([x,y,z,t]),NewSparseMultivariatePolynomial(Int
--R  eger,OrderedVariableList([x,y,z,t])),RegularTriangularSet(Integer,IndexedExpo
--R  nents(OrderedVariableList([x,y,z,t])),OrderedVariableList([x,y,z,t]),NewSpars
--R  eMultivariatePolynomial(Integer,OrderedVariableList([x,y,z,t]))))
--R                                                                 Type: Domain
--E 22
--S 23 of 23
toseSquareFreePart(pol,tower)$pack
--R
--R
--R                       2          3      5
--R   (23)  [[val= t y + z ,tower= {t  - 1,z  - t}]]
--IType: List(Record(val: NewSparseMultivariatePolynomial(...
--E 23
)spool
)lisp (bye)

———-

References^{[1] nothing}

8